OpenStudy (cookiimonster627):

find three additional polar representations. (-4,-5pi/4)

OpenStudy (arthur326):

I'm assuming that's just a point in the xy-plane described in polar coordinates. You have a lot of choices here (infinitely many, in fact). One option is to add any multiple of 2pi to the argument; that is \[(-4,-5\pi/4 + 2\pi k)\] is an equivalent representation for any integer k, since a full rotation about the origin doesn't change anything. Another thing you can do is use a positive radius instead, and offset the angle by pi: this works because a 180 degree rotation is the same as negating the radius, so if you do both, you get back the original point. So \[(4, -\pi/4 + 2\pi k)\] also describes the same point. Does that make sense?